# 4D Millimeter-Wave Radar in Autonomous Driving: A Survey

### v3

@@ VIII Conclusion
This paper offers a comprehensive overview of the role and potential of 4D mmWave radars in autonomous driving. It sequentially delves into the signal processing theory, datasets, learning-based data quality improvement methods, application algorithms in perception, localization and mapping. Furthermore, it casts a forward-looking gaze towards future trends and potential avenues for innovation in this rapidly evolving field. The exploration of 4D mmWave radars within the realm of autonomous driving is an ongoing endeavor. This comprehensive review serves as both a primer for those new to the field and a resource for seasoned researchers, offering insights into the current state of the art and highlighting the potential for future developments.

### v4

@@ VIII Conclusion
This paper offers a comprehensive overview of the role and potential of 4D mmWave radars in autonomous driving. It sequentially delves into the background theory, learning-based data generation methods, application algorithms in perception and SLAM, and related datasets. Furthermore, it casts a forward-looking gaze towards future trends and potential avenues for innovation in this rapidly evolving field. The exploration of 4D mmWave radars within the scope of autonomous driving is an ongoing endeavor. This comprehensive review serves as both a primer for those new to the field and a resource for experienced researchers, offering insights into the current state of the art and highlighting the potential for future developments.